OpenStudy (latinaas):

can anyone help me graph these three equations x = 2t y = t + 5, -2 ≤ t ≤ 3 x = 2t – 1 y = t2 + 5, -4 ≤ t ≤ 4 x = 6 sin t y = 6 cos t, 0 ≤ t ≤ 2π

OpenStudy (anonymous):

when t=-2 x=2*-2=-4 y=-2+5=3 point is (-4,3) when t=3 x=2*3=6 y=3+5=8 point is (6,8) again \[t=\frac{ x }{ 2 },y=\frac{ x }{ 2 }+5,2y=x+10,x-2y=-10\] it represents a straight line . so draw a line with above two points.

OpenStudy (latinaas):

thank you @surjithayer